A convex lens of focal length 50cm and a concave lens of focal length 100cm are separated by a distance of 10cm. What is the power of the
combination?

Grade:12th Pass

1 Answers

Chetan Mandayam Nayakar
11 years ago

after parallel light is incident on convex lens, object distance for concave lens= -40cm

(-1/40)+1/v = -1/100

1/v = 3/200

therefore, focal length=200/3 cm,power=3/200 Dioptre